# Learning to Discover Generalized Facial Expressions

### v1

@@ Abstract
We introduce Facial Expression Category Discovery (FECD), a novel task in the domain of open-world facial expression recognition (O-FER). While Generalized Category Discovery (GCD) has been explored in natural image datasets, applying it to facial expressions presents unique challenges. Specifically, we identify two key biases to better understand these challenges: Theoretical Bias—arising from the introduction of new categories in unlabeled training data, and Practical Bias—stemming from the imbalanced and fine-grained nature of facial expression data. To address these challenges, we propose FER-GCD, an adversarial approach that integrates both implicit and explicit debiasing components. In the implicit debiasing process, we devise F-discrepancy, a novel metric used to estimate the upper bound of Theoretical Bias, helping the model minimize this upper bound through adversarial training. The explicit debiasing process further optimizes the feature generator and classifier to reduce Practical Bias. Extensive experiments on GCD-based FER datasets demonstrate that our FER-GCD framework significantly improves accuracy on both old and new categories, achieving an average improvement of 9.8% over the baseline and outperforming state-of-the-art methods.
@@ 1 Introduction
Facial expression recognition (FER) is crucial in human-computer interaction, as expressions are a significant manifestation of human emotions [1, 2]. Traditional models are trained on seven basic facial expressions, i.e., happy, sad, surprise, anger, fear, disgust, and neutral. However, research indicates that humans can exhibit expressions that extend beyond these basic categories, including composite and complex expressions [3, 4], such as happy-surprise and perplexity. Consequently, models trained solely on these basic expressions perform poorly when encountering new types of expressions in real open-world FER (O-FER) scenarios. Additionally, manually annotating each expression in O-FER scenarios is prohibitively expensive and impractical. Consequently, reducing manual annotations and discovering generalized expression categories in O-FER have become important research focal points.
Recently, Open-Set FER [5, 6] has better solved the problem of models facing new expressions and is able to maintain high closed-set accuracy. However, current Open-Set FER methods still have limitations as they only aim at detecting new expressions without going for further new expression classification.
To further explore open-world scenarios and address previous limitations in O-FER, we introduced Generalized Category Discovery (GCD) [7] to FER scenarios for the first time, aiming to simultaneously recognize known expressions and discover different unknown expressions. Previous GCD approaches have made significant progress on general image datasets through research into contrastive learning techniques [7, 8, 9] and parametric classifier learning [10, 11].
However, directly applying existing GCD methods does not fully address the challenges outlined in this paper. These methods face inherent conflicts between recognizing previously learned categories and discovering new ones, resulting in biased learning. As training progresses, models tend to focus more on the new categories, causing a significant drop in accuracy for the old ones. Furthermore, the imbalanced distribution of expression data and the fine-grained differences between facial expressions exacerbate the bias, making it difficult for models to distinguish between subtle emotional variations. We summarize these challenges into two key points:
Figure 1: Left: GCD in O-FER aims to discover unkown (new), subtler facial expressions beyond the initial set of known (old) classes. Right: In GCD tasks, a major bias arises from new classes in unlabeled data interfering with model learning. In O-FER scenarios, this bias accumulates, sharply dropping old-class recognition accuracy in the mid-term. Our approach mitigates such biases and sustains a more stable improvement.
Challenge I: Theoretical Bias. In the semi-supervised GCD model, the introduction of unlabeled data with new emotion categories poses a challenge for the model’s learning, resulting in Theoretical Bias. During loss optimization, competition between old and new categories can lead to training bias, which is difficult to decouple in traditional GCD models with a single shared classification head, as shown in Fig. 9. Initially, the model improves in recognizing old categories due to its focus on labeled data. However, as training progresses, it increasingly prioritizes new categories, causing a significant drop in accuracy for the old categories, as shown in 1. This bias can be constrained through effective model design and optimization.
Challenge II: Practical Bias. In fine-grained and imbalanced O-FER scenarios, Practical Bias accumulates due to the characteristics of the data. Similarities between distinct expression categories and variations within the same category lead to misrecognition. The imbalanced data distribution favors majority expressions, resulting in lower accuracy for minority expressions. These accumulated biases hinder the model’s ability to capture distinguishing features, creating vague boundaries and difficulties in forming meaningful clusters.
Figure 2: Overview of FER-GCD framework. (i) In the implicit debiasing phase, we construct an adversarial process ($\mathcal{L}_{ad}$ ) between the Aux Head and the Main Head to help estimate the upper bound of bias and minimize it. (ii) In the explicit debiasing phase, we further enhance the performance of the feature generator $\psi$ and the Main Head through a combination of loss functions ($\mathcal{L}_{\text{rep}}$, $\mathcal{L}_{\text{cls}}$, $\mathcal{L}_{\text{bal}}$, $\mathcal{L}_{\text{cluster}}$). Clustering algorithm $\mathcal{L}_{\text{cluster}}$ will be introduced after the warm-up epochs.
In this paper, to address the challenges of Theoretical Bias and Practical Bias in O-FER, we propose a novel adversarial debiasing framework, named FER-GCD. FER-GCD integrates both implicit and explicit debiasing processes. The implicit debiasing process focuses on identifying and reducing the upper bound of Theoretical Bias, particularly caused by the introduction of new emotion categories. To achieve this, we define the F-discrepancy, which measures and minimizes the maximum bias in new categories. The explicit debiasing process addresses Practical Bias by enhancing the feature generator and classifier, leading to more discriminative feature boundaries that mitigate the effects of data imbalance and fine-grained variations. Together, these debiasing techniques significantly improve the robustness and accuracy of FER-GCD in tackling both types of bias in open-world FER scenarios.
Our contributions can be summarized as follows: (i) We introduce Facial Expression Category Discovery, aiming to understand human emotions in real and various scenarios. To the best of our knowledge, this is the first time to address this issue. (ii) We identify two main biases: Theoretical Bias, caused by new categories in unlabeled training data, and Practical Bias, arising from fine-grained and imbalanced facial expression scenarios. (iii) We propose FER-GCD, a novel adversarial framework integrating implicit and explicit debiasing techniques. (iv) Extensive experiments conducted on GCD-based FER datasets demonstrate approach’s superiority over other state-of-the-art GCD methods.
2 FER-GCD algorithm
In this section, we introduce the debiasing strategy of FER-GCD, which addresses both Theoretical Bias and Practical Bias through implicit and explicit debiasing. Specifically, in section 2.2, we define and constrain Theoretical Bias through mathematical bounds, followed by an adversarial training process for implicit debiasing. In section 2.3, we tackle Practical Bias through explicit debiasing, which enhances the feature generator and the main classification head to handle imbalanced and fine-grained facial expression data.
2.1 Preliminaries
@@
Problem Setting. The GCD [7] problem involves maintaining the model’s ability to categorize known classes while also discovering new classes when given partially labeled data and a large amount of unlabeled data. In the context of FER datasets, we define the unlabeled dataset as $\mathcal{D}_{U}={(\mathbf{x}_{i}^{u},\mathbf{y}_{i}^{u})}\in\mathcal{X}\times\mathcal{Y}_{u}$, where $\mathcal{Y}_{u}$ is the label space for unlabeled data points. The goal of GCD in this context is to train a model that effectively categorizes the instances in $\mathcal{D}_{U}$ using information from a labeled dataset $\mathcal{D}_{L}={(\mathbf{x}_{i}^{l},\mathbf{y}_{i}^{l})}\in\mathcal{X}\times\mathcal{Y}_{l}$, where $\mathcal{Y}_{l}$ is the label space for labeled data points, and $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$. In this study, we assume the number of categories in the unlabeled space is known and is $\mathcal{Y}_{u}$, represented by $K_{u}=|\mathcal{Y}_{u}|$.
@@
2.2 Implicit Debiasing for Theoretical Bias
Implicit debiasing targets Theoretical Bias, which arises from new categories in unlabeled data. We first estimate and constrain this bias, then minimize its upper bound through adversarial training.
Definition 1 (Metrics of bias)
With a sufficiently large amount of data, we define $\xi(\cdot,\cdot)$ as a measure of the difference between the model’s predictions and the ground truth, where the predictions go through a softmax layer, mapping to the probability space $\mathbb{P}$. A mapping function $f:\mathbb{P}\rightarrow\mathbb{E}$ then converts these probabilities into the Euclidean space $\mathbb{E}$. We can define the metric:
The non-negativity, symmetry and metricity satisfied by $\xi(\cdot,\cdot)$ will help us to carry out the proofs that follow, please see the appendix 1 for the specific properties.
Since GCD models all use the semi-supervised training strategy, we define the prediction discrepancy of the models on labeled data $\mathcal{D}_{L}$ and unlabeled data $\mathcal{D}_{U}$ separately, we use the metric defined in defination1:
$\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})=\mathbb{E}_{x\in\mathcal{D}_{L}}\left[\xi\left(\mathcal{H}(x),\mathcal{F}(x)\right)\right],\quad\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})=\mathbb{E}_{x\in\mathcal{D}_{U}}\left[\xi\left(\mathcal{H}(x),\hat{\mathcal{F}}(x)\right)\right],$ (2)
Definition 2 (Metric promotion based on category space)
In GCD task, $\mathcal{D}_{U}$ contains both new and old categories that are inaccessible to us. The proportion of old and new classes can be assumed to follow a binomial distribution, where the probability of old class is denoted as $\theta$. Therefore, we can concretely define that the discrepancy on unlabeled data consists of the following convex combination:
$\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})=(1-\theta)\xi^{new}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})+\theta\xi^{old}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}}),$ (3)
Furthermore, since $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$, and based on empirical extrapolations, we can assume that the model will make more mistakes on unlabeled datasets (even if the categories are the same), mathematically represented as follows:
@@ 2.2 Implicit Debiasing for Theoretical Bias
We define the upper bound on the discrepancy between the current model $\mathcal{H}$ and an arbitrary model hypothesis $\mathcal{H^{\prime}}$ on both labeled and unlabeled data. The F-discrepancy is:
where $\alpha$ is a tuning parameter that adjusts the weight of the prediction bias for labeled data. The F-discrepancy metric is crucial for the subsequent lemma 1.
Lemma 1 (Upper Bound on Theoretical Bias from New Categories)
Since we cannot access the labels of $\mathcal{D}_{U}$ to determine whether the data belongs to new categories or old categories, we need to leverage $\mathcal{D}_{L}$ to help constrain the estimation of the upper bound of the new category bias in a rigorous manner. Eventually we get an upper bound on the bias of the new category:
$\xi^{new}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\frac{1}{1-\theta}\left[(\alpha-\theta)\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})+\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})+\lambda\right],$ (6)
Where $\lambda=\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H^{*}},\mathcal{F})+\xi_{\mathcal{D}_{U}}(\mathcal{H^{*}},\mathcal{F})$. Proof 2 is provided in the appendix.
Adversarial Optimization for Theoretical Bias: In Lemma 1, we establish an upper bound for the Theoretical Bias introduced by new categories in the unlabeled data. The goal of the implicit debiasing process is to minimize this upper bound. This serves to constrain the learning of new categories and helps avoid the pitfall of degrading the performance of previously learned categories due to the introduction of new data. This min-max process is computed as follows:
@@
$\mathop{\min}_{\mathcal{H}_{h,\psi}}\mathop{\max}_{\mathcal{H^{\prime}}_{h_{a},\psi_{a}}}(\alpha-\theta)\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})+\left|\xi_{\mathcal{D}_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|,$ (7)

### v2

# DIG-FACE: De-biased Learning for Generalized Facial Expression Category Discovery

@@ Abstract
We introduce a novel task, Generalized FAcial expression Category discovEry (G-FACE), that discovers new, unseen facial expressions while recognizing known categories effectively. Even though there are generalized category discovery methods for natural images, they show compromised performance on G-FACE. We identified two biases that affect the learning: implicit bias, coming from an underlying distributional gap between new categories in unlabeled data and known categories in labeled data, and explicit bias, coming from shifted preference on explicit visual facial change characteristics from known expressions to unknown expressions. By addressing the challenges caused by both biases, we propose a Debiased G-FACE method, namely DIG-FACE, that facilitates the de-biasing of both implicit and explicit biases. In the implicit debiasing process of DIG-FACE, we devise a novel learning strategy that aims at estimating and minimizing the upper bound of implicit bias. In the explicit debiasing process, we optimize the model’s ability to handle nuanced visual facial expression data by introducing a hierarchical category-discrimination refinement strategy: sample-level, triplet-level, and distribution-level optimizations. Extensive experiments demonstrate that our DIG-FACE significantly enhances recognition accuracy for both known and new categories, setting a first-of-its-kind standard for the task.
@@ 1 Introduction
Facial expression recognition (FER) is important in various tasks like the human-computer interaction, as expressions are a significant manifestation of human emotions [20, 18]. Traditional models are trained to recognize seven basic facial expressions, i.e., happy, sad, surprise, anger, fear, disgust, and neutral. However, researches [16, 6, 23] indicate that humans can exhibit expressions that extend beyond these basic categories, including composite and complex expressions, such as happy-surprise and perplexity. Sometimes, human expressions contain many complex and unexplored subtleties and variations [49]. When encountering new types of expressions in real-world scenarios, models trained solely on basic expressions would perform poorly. Furthermore, manually annotating each expression in open-world FER scenarios could be prohibitively expensive and impractical. Therefore, reducing manual annotations and discovering generalized facial expression categories have become increasingly necessary when exploring the open-set setting.
Figure 1: G-FACE aims to discover unknown (new) facial expressions while recognizing known (old) classes, both of which are present in the unlabeled data.
Recent advancements in open-set FER [45, 24] have improved the ability of models to detect new, unknown expressions while maintaining high accuracy on known categories. However, these methods primarily focus on identifying one unknown category alongside known categories, limiting their effectiveness in scenarios with multiple unknown expressions and hindering their applicability in real-world contexts. Then, the need for addressing this limitation has driven the development of Generalized Category Discovery (GCD) [35]. The GCD employs semi-supervised training to utilize both labeled and unlabeled data, assigning cluster IDs to the unlabeled data to enable recognition of both known and multiple unknown categories. Although GCD has shown success on natural image datasets [37, 26], we found that directly applying it to facial expression recognition (FER) introduces unique challenges. Unlike general images, FER data involves much more imbalanced distributions of known and unknown categories, as well as much more subtle variations between expressions, making it extremely difficult for current GCD methods to maintain accuracy on known categories while identifying new ones. In this study, to specifically formulate the GCD task for FER, we draw inspiration from the typical GCD formulation and introduce a novel task, namely Generalized FAcial expression Category discovEry (G-FACE), to represent the task that simultaneous identifies known facial expressions and discovers new, unseen facial expressions in open-set data.
As mentioned above, the G-FACE presents unique challenges, which we could conclude as a biasing problem. More specifically, we observe implicit and explicit biases that would set obstacles to the G-FACE:
$\bullet$ Challenge I: Implicit Bias. During the semi-supervised learning for GCD, unlabeled data has new categories whose effective decision boundaries usually subtly shifted from known categories. This introduces implicit distributional gaps between known and unknown data, which increases the learning difficulties for the G-FACE. Fig. 2 (top) partially illustrates the effect caused by the implicit bias, where this shift leads to a significant drop in accuracy for known expression categories as training goes on (e.g. possibly the effective decision boundaries of known have been shifted).
$\bullet$ Challenge II: Explicit Bias. We identify explicit bias as the shifts in the preference of visible subtle facial characteristics from known expressions to unknown expressions. Different from implicit bias representing shifted decision boundaries, the differences in key facial characteristics between known and unknown expressions would make learned decision boundaries unclear and less distinguishable for both known and unknown categories. This can be partially supported by overlapped known/unknown data distributions, as shown in Fig. 2 (bottom) .
Figure 2: Top: Implicit bias leads to a decline in recognition accuracy for known categories in later training stages, as seen in previous GCD models like Baseline SimGCD. Bottom: Explicit bias can be analyzed by observing the overlap between known and unknown classes, which may lead to a blurred decision boundary. We present a t-SNE visualization of the baseline SimGCD.
To address these issues, we propose a novel De-bIased G-FACE framework, DIG-FACE, which incorporates both implicit and explicit debiasing strategies. In the implicit debiasing step, we introduce a novel F-discrepancy-based metric to help reveal and reduce the upper bound of implicit bias. The explicit debiasing step enhances visual feature learning through a three-level category-discrimination refinement learning strategy to address fine-grained subtle distinctions. Together, these strategies significantly improve the model’s robustness and accuracy in the proposed G-FACE task. Our contributions can be summarized as:
(i) We propose a novel G-FACE task, which aims to simultaneously recognize known facial expressions and discover unknown categories. To the best of our knowledge, this is the first time this problem has been formulated in the FER literature.
(ii) We identify two core biases that introduce challenges in G-FACE: implicit bias from underlying distributional gap, and explicit bias from diversified preference on subtle visual facial characteristics, which may confuse decision boundaries of known/unknown categories, complicating accurate recognition.
(iii) We propose DIG-FACE, incorporating both implicit and explicit debiasing processes to address the identified biases. The DIG-FACE has shown outstanding benefits on the G-FACE task, consistently outperforming state-of-the-art GCD approaches on various datasets by a great margin.
2 Related Work
Facial Expression Recognition (FER) conveys rich information and is a key focus of current AI research [8, 33, 31, 3, 14, 20]. There have also been many studies in recent years aimed at improving the performance of FER models [41, 38, 19, 44]. For instance, Yang et al[41] introduce a de-expression learning method. Wang and his team propose an effective Self-Cure Network (SCN). Besides, semi-supervised learning [19] and challenging imbalance problems [44] are also important research directions in FER. Recently, in order to further explore open-world settings, Zhang et al[45] introduce open-set setting to FER scenario. In addition, Liu et al. proposed video-baset open-set setting in FER [24]. However the open-set task aims to detect new classes that do not belong to the previous known category without further categorization, which remains a limitation.
Category Discovery aims to discover new categories from unlabeled data. During training, labeled and unlabeled training data provide important visual conceptual information. The assumption of Novel Category Discovery (NCD) [10, 13, 12, 46] is that there is no overlap in categories between the labeled and unlabeled datasets. However, this assumption is unrealistic in the broader open-world [35, 29, 43] setting, where the unlabeled dataset not only encompasses categories previously learned by the model from labeled data but also novel categories. Prior studies have devised efficient techniques for category discovery, utilizing parametric classifiers [39, 26], enhancing representation learning [29, 9], or employing prompting learning with larger models [43]. In addition, some researches aim to focus on more realistic or difficult GCD problems, such as incremental learning-GCD [47], long-tail recognition GCD [22], cross-domain GCD [36], etc. However, most existing GCD methods utilize a single shared classification head, thus failing to decouple the accumulation of biases.
3 Problem Formulation
In this section, we formalize the task definition of G-FACE and provide a detailed explanation of its two major challenges: Implicit Bias and Explicit Bias. Tab. 1 provides the preliminary notations.
3.1 Problem Setting
@@
Following GCD setting [35], the G-FACE problem focuses on preserving the model’s capability to categorize known expression classes and simultaneously discover new ones from partially labeled and abundant unlabeled data. We define the unlabeled FER dataset as $\mathcal{D}_{U}=\{(\mathbf{x}_{i}^{u},\mathbf{y}_{i}^{u})\}\subset\mathcal{X}\times\mathcal{Y}_{u}$, where $\mathcal{Y}_{u}$ is the label space for unlabeled data points. The goal of G-FACE is to train a model that effectively categorizes the instances in $\mathcal{D}_{U}$ using information from a labeled FER dataset $\mathcal{D}_{L}=\{(\mathbf{x}_{i}^{l},\mathbf{y}_{i}^{l})\}\subset\mathcal{X}\times\mathcal{Y}_{l}$, where $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$. In this study, we adopt a parametric approach [39, 26, 1] and assume the number of categories in the unlabeled space is known, denoted as $K_{u}=|\mathcal{Y}_{u}|$. The evaluation metric is defined in 5.1.
3.2 Implicit Bias
We identify the implicit bias which subtly shifts decision boundaries and creates distributional gaps. While this bias cannot be directly observed from the image data, it can be constrained through mathematical techniques. To address it, we define a discrepancy metric to measure the divergence between predictions and true labels and decompose the overall discrepancy to isolate the contribution of new categories. We then introduce the F-discrepancy, a theoretical tool that characterizes the relationship between discrepancies on labeled and unlabeled data. Based on this, we derive a formal upper bound for implicit bias. The key notions are summarized in Tab 1.
Definition 1 (Discrepancy Metric)
To quantify the discrepancy between model predictions and true labels, we define a discrepancy metric $\xi(\cdot,\cdot)$. For a set of data points $\{x_{i}\}_{i=1}^{n}$, the model predictions are obtained through a softmax function and then mapped to Euclidean space $\mathbb{E}$ using a function $f:\mathbb{P}\rightarrow\mathbb{E}$:
$\footnotesize\xi(\mathcal{H}_{h,\psi}(x),\mathcal{F}(x))=\sqrt{\sum_{i=1}^{n}\left\|\mathcal{H}(x_{i})-\mathcal{F}(x_{i})\right\|^{2}},$ (1)
where $\mathcal{H}$ denotes the model and $\mathcal{F}$ represents the true label function. This metric is used to calculate discrepancies on labeled data $\mathcal{D}_{L}$ and unlabeled data $\mathcal{D}_{U}$ separately:
$\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})=\mathbb{E}_{x\in\mathcal{D}_{L}}\left[\xi\left(\mathcal{H}(x),\mathcal{F}(x)\right)\right],$ (2)
$\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})=\mathbb{E}_{x\in\mathcal{D}_{U}}\left[\xi\left(\mathcal{H}(x),\hat{\mathcal{F}}(x)\right)\right],$
where $\mathcal{F}$ and $\hat{\mathcal{F}}$ denote the labeling functions for $\mathcal{D}_{L}$ and $\mathcal{D}_{U}$, respectively.
Table 1: List of symbols for DIG-FACE.
Symbol Description
$\mathbb{R}$ The model hypothesis space
$\mathcal{D}_{L}$ Labeled dataset with $\mathcal{N}$ known categories
$\mathcal{D}_{U}=\{\mathcal{D}_{U}^{\text{old}},\mathcal{D}_{U}^{\text{new}}\}$ Unlabeled dataset with $\mathcal{D}_{U}^{\text{old}}$ (known, $\mathcal{N}$ categories) and $\mathcal{D}_{U}^{\text{new}}$ (new, $\mathcal{M}$ categories)
@@ 3.2 Implicit Bias
$\mathcal{H}$ or $\mathcal{H}_{h,\psi}\in\mathbb{R}$ Main model $\mathcal{H}$ consists of feature exactor $\psi$ and classification head $h$
$\mathcal{H}_{h^{*},\psi^{*}}\in\mathbb{R}$, $\mathcal{H^{\prime}}_{h_{a},\psi_{a}}\in\mathbb{R}$ Optimal reference model; auxiliary model for bounding bias
$\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})$, Discrepancy metrics on $\mathcal{D}_{L}$ and
$\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})$ $\mathcal{D}_{U}$
$\mathbb{P}$, $\mathbb{E}$ Probability space of model predictions; Euclidean space
$f:\mathbb{P}\rightarrow\mathbb{E}$ Mapping function from probability to Euclidean space
$\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})$ F-discrepancy
Figure 3: Overview of DIG-FACE framework. (i) In the implicit debiasing stage, we estimate and minimize the maximum bias by enforcing consistency between the main and auxiliary head on $\mathcal{D}_{L}$ and inconsistency on $\mathcal{D}_{U}$. At this stage, $P_{l}$ and $P_{u}$ denote the pseudo-labels of labeled and unlabeled data. (ii) In the explicit debiasing stage, we enhance recognition of known and unknown categories through sample-level, triplet-level, and distribution-level optimization, and improve decision-making under G-FACE with parametric learning.
Definition 2 (Category-based Discrepancy Composition)
In G-FACE, the unlabeled dataset $\mathcal{D}_{U}$ contains both new and old categories. Assuming the proportion of old classes follows a probability $\theta$, the discrepancy on $\mathcal{D}_{U}$ can be decomposed as:
$\footnotesize\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})=(1-\theta)\xi^{new}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})+\theta\xi^{old}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}}),$ (3)
where $\xi^{new}_{\mathcal{D}_{U}}$ and $\xi^{old}_{\mathcal{D}_{U}}$ denote discrepancies for new and old categories, respectively. Given $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$, we assume that the model’s error on $\mathcal{D}_{U}$ is at least as large as on $\mathcal{D}_{L}$:
$\footnotesize\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})\leq\xi^{old}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\xi_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}}).$ (4)
@@
The F-discrepancy is defined as an upper bound on the divergence between discrepancies of $\mathcal{D}_{L}$ and $\mathcal{D}_{U}$. It compares the current model $\mathcal{H}$ with an auxiliary hypothesis $\mathcal{H^{\prime}}$ across the hypothesis space $\mathcal{R}$, and is expressed as:
@@
$\footnotesize\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})=\sup_{\mathcal{H},\mathcal{H^{\prime}}\in\mathcal{R}}\left|\xi_{\mathcal{D}_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|,$ (5)
where $\alpha$ adjusts the contribution of labeled data discrepancies. This definition provides a foundation for bounding implicit bias in subsequent analysis.
Lemma 1 (Upper Bound on Implicit Bias)
As the labels of $\mathcal{D}_{U}$ are not directly observable, we leverage the labeled data $\mathcal{D}_{L}$ to establish an upper bound for the implicit bias caused by new categories. Specifically, the implicit bias satisfies:
$\footnotesize\xi^{new}_{\mathcal{D}_{U}}(\mathcal{H},\hat{\mathcal{F}})\leq\frac{1}{1-\theta}\left[(\alpha-\theta)\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{F})+\mathop{\Delta}(\mathcal{D}_{U},\mathcal{D}_{L})+\lambda\right],$ (6)
where $\lambda=\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H^{*}},\mathcal{F})+\xi_{\mathcal{D}_{U}}(\mathcal{H^{*}},\mathcal{F})$. The proof is provided in appendix.
3.3 Explicit Bias
Compared to implicit bias, explicit bias is more easily observed. It typically manifests as subtle visual features that cause blurry decision boundaries and difficulties in distinguishing categories. While past FER methods have made significant progress in addressing closed-set [42, 44] and open-set FER [45, 24] that do not require further classification of new classes, we identify that G-FACE faces a unique challenge: the subtle differences in the preference on different visual features between known and unknown expressions produce overlapped distributions in the feature spaces, making it difficult to distinguish diverse categories. This phenomenon can be partially reflected in the t-SNE visualization of the baseline SimGCD (Fig. 2, bottom), where the overlap in data distributions indicates a clear bias.
4 DIG-FACE Debiasing Methodology
4.1 Implicit Debiasing
In Eq. (6) of Lemma 1, we provide an upper bound for the implicit bias, and thus our goal is to minimize this upper bound to control the learning of new classes. Specifically, it requires us to leverage the inconsistency between the auxiliary head $\mathcal{H}^{\prime}$ and the main head $\mathcal{H}$ on $\mathcal{D}_{U}$, while ensuring their consistency on $\mathcal{D}_{L}$. Intuitively, this will enable the feature exactor to maintain robust performance even under poorer decision boundaries, thereby effectively mitigating the impact of decision boundary shifts caused by the implicit bias. The Min-Max process is as follows:
$\footnotesize\mathop{\min}_{\mathcal{H}}\mathop{\max}_{\mathcal{H^{\prime}}}(\alpha-\theta)\cdot\xi_{D_{L}}(\mathcal{H},\mathcal{F})+\left|\xi_{\mathcal{D}_{U}}(\mathcal{H},\mathcal{H^{\prime}})-\alpha\cdot\xi_{\mathcal{D}_{L}}(\mathcal{H},\mathcal{H^{\prime}})\right|,$ (7)